OpenStudy (anonymous):

Find f-1(x) (-1 is an inverse function). Sketch the graphs of y=f(x) and y=f-1(x) on the same coordinate axes. Answer for f(x)=x^3+1

OpenStudy (anonymous):

f(x) = x^3 + 1 y = x^3 +1 x = y^3 +1 x-1 = y^3 cube root) of (x-1) = y plug in values of x in the function and in the inverse